# Directional anomaly detection

### v1

@@ 4 Experiments on synthetic datasets
ai4i2020 [15] 10 000 Simulated machine operation records 339 Five different failure modes (not distinguished by the label) 6
diabetes-risk [8] 520 Patients with diabetes symptoms 320 Patients actually diagnosed with diabetes 14
@@
heart-failure [1] 299 Heart patients 96 Patients who died in the follow-up period 11
@@
south-german-credit [6] 1000 Credits provided by a bank 300 Bad credits 20
@@
wpbc [20] 138 Patients who have undergone surgery for an invasive malignant breast cancer tumor with no evidence of distant metastases 28 Patients who experienced recurrence within two years 32
Table 3: Mean cross-validation AUROC for the absolute, ramp and signed distance variants. Bold: highest value (before rounding) for NND and ALP, respectively.
Dataset NND ALP
Absolute Ramp Signed Absolute Ramp
ai4i2020 0.823 0.922 0.724 0.877 0.924
diabetes-risk 0.971 0.923 0.716 0.895 0.926
fertility 0.602 0.653 0.540 0.581 0.636
heart-failure 0.715 0.769 0.735 0.734 0.766
phishing-websites 0.901 0.927 0.804 0.927 0.936
post-operative 0.476 0.504 0.557 0.459 0.484
qualitative-bankruptcy 1.000 1.000 0.998 1.000 1.000
south-german-credit 0.648 0.718 0.683 0.648 0.714
thoraric-surgery 0.597 0.624 0.583 0.634 0.621
wdbc 0.950 0.976 0.969 0.957 0.981
wisconsin 0.995 0.994 0.995 0.872 0.995
wpbc 0.570 0.625 0.633 0.537 0.654
5 Experiments on real-life datasets
@@
We will now evaluate our proposal on a number of real-life datasets from the UCI repository [3] that can be approached as directional anomaly detection problems. The main properties of these datasets are listed in Table 2. The attributes of these datasets express machine operating conditions (ai4i2020), medical symptoms, comorbidities and lifestyle factors (diabetes-risk, fertility, heart-failure, post-operative and thoraric-surgery), tumor characteristics (wdbc, wisconsin and wpbc), and risk indicators provided by experts (phishing-websites, qualitative-bankrupty and south-german-credit). ai4i2020 and post-operative contain a mixture of directional and adirectional attributes, the other datasets only have directional attributes.
As in the previous section, we use AUROC to evaluate anomaly detection performance. We perform 5-fold cross-validation on the normal records from each dataset, creating at each iteration a test set by combining one fifth of the normal records with all anomalous records. For NND, we use linear weights and fix $k=8$. We rescale all attributes by subtracting the midhinge and dividing by the semi-interquartile range of the normal training values, such that the interquartile range becomes $[-1,1]$ in the training data.
Table 3 lists the mean cross-validation AUROC obtained by the various variants for each dataset. The first thing to note is that for NND, unlike what we saw for synthetic data in the previous section, ramp distance generally performs (much) better than signed distance. The only datasets for which signed distance performs better are post-operative and wpbc, for which the AUROC scores are very low anyway. When we perform a one-side Wilcoxon signed-rank test, we obtain $p=0.021$. We can conclude from this that real-life anomaly detection problems are different from our synthetic datasets in one important aspect: an unexpectedly low score on one risk factor does not compensate for high scores on other risk factors.
However, these experiments do confirm the usefulness of directional anomaly detection, because ramp distance performs about as well or better than absolute distance. For NND, the signed-rank test gives $p=0.011$, or $p=0.023$ if we apply Holm-Bonferroni correction [7] for multiple testing, while for ALP it gives $p=0.0029$.
@@
There is one notable exception: diabetes-risk for NND. A potential explanation for this is the fact that the attributes of this dataset encode diabetes symptoms, and that moreover, the normal records do not represent average healthy people, but non-diabetes patients who nevertheless display symptoms of diabetes. Consequently, this may undermine the directionality of this dataset — diabetes patients may not necessarily have more symptoms than this particular group of other patients, but rather different combinations of symptoms, which would make absolute distance a better fit than ramp distance.
@@
Indeed, when we look at the mean attribute values of normal and anomalous records in this dataset, we find that for five of the fourteen attributes, the anomalous records (diabetes patients) have lower or only slightly higher mean values than the normal records. If we rerun the experiment while treating these five attributes as non-directional, the AUROC obtained by ramp distance matches that of absolute distance.
@@ 6 Conclusion
In this paper, we have introduced directional anomaly detection, a new problem setting wherein only high (or only low) values of certain attributes are indicative of anomality, and which can therefore be viewed as risk factors. We have proposed two different ways of adapting distance-based anomaly detection algorithms to make use of this knowledge: ramp distance and signed distance. The difference between these two variants corresponds to a choice in how a practitioner wants to interpret an unexpectedly low value on a given risk factor. With ramp distance, such low values are simply discounted, whereas with signed distance, they contribute negative evidence against anomality. Both variants contrast with absolute distance, the non-directional baseline wherein such low values contribute positive evidence for anomality.
In an experiment with synthetically generated Gaussian and Bernoulli data, we found that signed distance is able to achieve slightly higher anomaly detection than ramp distance. However, in a subsequent experiment with real-life datasets, ramp distance performed substantially better. We conclude from this that in typical real-life use cases, unexpectedly low values on risk factors should not compensate for high values on other risk factors.
We also found that ramp distance generally performed as good or better than absolute distance. Therefore, we recommend the use of ramp distance when a practitioner knows that some of the attributes in their dataset are really risk factors.
Finally, we will end by briefly considering what it means if, for a given dataset, ramp distance should nevertheless result in worse anomaly detection performance than absolute distance. One possibility is that the directional hypothesis is false for one or more of the attributes, that there is a good reason after all why (some) anomalies should have lower values than normal. This is in fact what we saw with one of the real-life datasets that we evaluated, the only dataset where absolute distance clearly performed better than ramp distance for NND.
The other possibility is that there is no clear explanation, no plausible causal relation linking a lower attribute value to a higher risk of abnormality. In that case using ramp distance may still be preferable, because it better matches the domain knowledge of the practitioner, resulting in a more interpretable prediction model.
Acknowledgements

### v2

@@ 5.2 Synthetic datasets
ai4i2020 [15] 10 000 Simulated machine operation records 339 Five different failure modes 6
diabetes-risk [8] 520 Patients with diabetes symptoms 320 Actual diabetes patients 14
@@
heart-failure [1] 299 Heart patients 96 Patients who died 11
@@
south-german-credit [7] 1000 Bank credits 300 Bad credits 20
@@
wpbc [20] 138 Patients with surgery for invasive malignant breast cancer 28 Patients who experienced recurrence within two years 32
5.3 Real-life datasets
@@
We have selected 10 real-life datasets from the UCI repository [4] that can be approached as monotonic anomaly detection problems. The main properties of these datasets are listed in Table 2. The attributes of these datasets express machine operating conditions (ai4i2020), medical symptoms, comorbidities and lifestyle factors (diabetes-risk, fertility, heart-failure, post-operative and thoraric-surgery), tumor characteristics (wdbc, wisconsin and wpbc), and risk indicators provided by experts (phishing-websites, qualitative-bankruptcy and south-german-credit). ai4i2020 and post-operative contain a mixture of monotonic and ordinary attributes, the other datasets only have monotonic attributes.
On these datasets, we perform 5-fold cross-validation on the normal records, creating at each iteration a test set by combining one fifth of the normal records with all anomalous records.
5.4 Hyperparameters
The anomaly detection algorithms from Section 3 have a number of hyperparameters which can in principle be tuned when there are anomalies available for validation purposes [13]. In general, however, this is not the case in the semi-supervised setting, so we will use sensible default values.
For NND and LOF, 1-distance (Boscovich distance) generally outperforms 2-distance (Euclidean distance) [12, 11], and for ALP, 1-distance is the standard choice [12]. For consistency, we also choose 1-distance for CD. For both weighted NND and ALP, the standard choice is to use linearly descending weights [12, 11]. As for the number of neighbours, $k=8$ has been established as the optimal default value for weighted NND [11], $k=2.5\log n$ for LOF [12] and $k=5.5\log n$ and $l=6\log n$ for ALP [12], where $n$ is the number of training records. For CD, we will use the marginal mean (centroid) as the centre value.
Finally, while not typically seen as a hyperparameter, all distance-based algorithms also depend on the relative scale of the attributes. We normalise the real-life datasets by dividing each attribute by the semi-interquartile range of the normal training values, such that the interquartile range becomes $[-1,1]$ in the training data.
Figure 1: Mean AUROC obtained with synthetically generated datasets with Gaussian and Bernoulli attributes. $a$: distance between the distributions of anomalous and normal attribute values; $b$: difference between the probability of a positive anomalous and a positive normal attribute value.
6 Results
6.1 Synthetic datasets
The results for the synthetic datasets are displayed in Figure 1.
With Gaussian attributes, the results are substantially the same for the different anomaly detection algorithms. Ramp distance clearly performs better than absolute distance, in line with our monotonic hypothesis. For CD, signed distance performs even better than ramp distance, but the difference is small.
With Bernoulli attributes, ramp distance also consistency performs well, but there is more variation for absolute distance. For ALP and NND, the performance deficit with respect to ramp distance is much smaller than with Gaussian attributes, and for CD all three distance measures give practically identical results. Finally, for LOF, the results with absolute distance are particularly poor.
Taken together, these results suggest that when we know for a given attribute that anomalies have a higher mean value than normal records, we should indeed use ramp distance rather than absolute distance. On these idealised datasets, CD performs as well as the nearest neighbour algorithms, and signed distance appears to produce even better results than ramp distance. However, we will see in the next subsection that this does not carry over to the more complicated structure of real-life datasets.
Table 3: Mean cross-validation AUROC for the absolute (A), ramp (R) and signed (S) distance variants. Bold: highest value (before rounding) for each algorithm.
Dataset CD NND LOF ALP
A R S A R A R A R
ai4i2020 0.807 0.868 0.853 0.823 0.922 0.867 0.909 0.877 0.924
diabetes-risk 0.886 0.778 0.715 0.971 0.923 0.745 0.920 0.895 0.926
fertility 0.388 0.478 0.540 0.602 0.653 0.534 0.605 0.581 0.636
heart-failure 0.714 0.773 0.735 0.715 0.769 0.736 0.715 0.734 0.766
phishing-websites 0.665 0.767 0.804 0.901 0.927 0.851 0.902 0.927 0.936
post-operative 0.481 0.522 0.560 0.476 0.504 0.522 0.527 0.459 0.484
qualitative-bankruptcy 1.000 1.000 0.998 1.000 1.000 1.000 1.000 1.000 1.000
south-german-credit 0.613 0.711 0.683 0.648 0.718 0.638 0.675 0.648 0.714
thoraric-surgery 0.574 0.629 0.583 0.597 0.624 0.607 0.542 0.634 0.621
wdbc 0.953 0.969 0.969 0.950 0.976 0.950 0.954 0.957 0.981
wisconsin 0.995 0.995 0.995 0.995 0.994 0.692 0.990 0.872 0.995
wpbc 0.529 0.624 0.633 0.570 0.625 0.555 0.582 0.537 0.654
6.2 Real-life datasets
Table 3 lists the mean cross-validation AUROC obtained by the various variants for each real-life dataset. CD does not perform quite as well as the other algorithms, which is not surprising given the more complicated structure of these datasets. However, it is remarkable that it performs as well as it does. Overall, NND and ALP achieve the best performance and consistently outperform LOF.
The main question that we want to answer is whether monotonic anomaly detection outperforms ordinary anomaly detection on monotonic datasets. Note first that contrary to what we saw for synthetic datasets, signed distance no longer outperforms ramp distance for CD. Moreover, it is consistently beaten by the combination of NND with ramp distance. This leaves ramp distance as the clearly preferred monotonic variant. In line with our monotonic hypothesis, we find that ramp distance significantly outperforms absolute distance: the one-sided Wilcoxon signed-rank test gives $p=0.030$ (CD), $p=0.011$ (NND), $p=0.025$ (LOF) and $p=0.0029$ (ALP).
@@ 6.2 Real-life datasets
In those cases where absolute distance performs better than ramp distance, the difference is very small or the AUROC scores are altogether very low. There is one notable exception: the diabetes-risk dataset for CD and NND. A potential explanation for this is the fact that the attributes of this dataset encode diabetes symptoms, and that moreover, the normal records do not represent average healthy people, but non-diabetes patients who nevertheless display symptoms of diabetes. Consequently, this may undermine the monotonicity of this dataset — diabetes patients may not necessarily have more symptoms than this particular group of other patients, but rather different combinations of symptoms, which would make absolute distance a better fit than ramp distance.
@@
Indeed, when we look at the mean attribute values of normal and anomalous records in this dataset, we find that for five of the fourteen attributes, the anomalous records (diabetes patients) have lower or only slightly higher mean values than the normal records. If we rerun the experiment while treating these five attributes as non-monotonic, the AUROC obtained by ramp distance matches that of absolute distance.
@@ 7 Conclusion
In this paper, we have introduced monotonic anomaly detection, a new problem setting that incorporates the domain knowledge that only high (or only low) values of certain attributes are indicative of anomality (and which can therefore be viewed as risk factors). We have shown that distance-based anomaly detectors can take this monotonicity into account by using an asymmetric distance function. For this purpose, we have proposed two variants of the commonly used family of Minkowski $p$-distances: ramp distance and signed distance. The difference between these two variants corresponds to a choice in how a practitioner wants to interpret a relatively low value of a given attribute. With ramp distance, such low values are simply discounted, whereas with signed distance, they contribute negative evidence against anomality. Both variants contrast with absolute distance, the non-monotonic baseline wherein such low values contribute positive evidence for anomality.
However, because signed distance lets low values of some attributes compensate for high values of other attributes, the precise distribution of the attribute values becomes irrelevant, and anomality is purely determined on the basis of the total attribute value sum. This makes signed distance fundamentally incompatible with nearest neighbour algorithms, which are based on the premise that the normal data class is characterised by a specific distribution of attribute values.
This contrast is reflected in our results on synthetic and real-life data. On the idealised synthetic datasets, the simple CD algorithm performed about as well as the nearest neighbour algorithms, and signed distance achieved slightly higher performance than ramp distance. However, on the more complex real-life datasets, the nearest neighbour algorithms with ramp distance outperformed CD with signed distance. Therefore, ramp distance is the preferred monotonic variant. When comparing ramp distance with absolute distance, which corresponds to ordinary anomaly detection, we found that on both synthetic and real-life datasets, ramp distance performs significantly better.
We conclude that practitioners faced with a monotonic anomaly detection problem — where they know that only higher values of certain attributes are indicative of anomality — should use a nearest neighbour algorithm that uses ramp distance for monotonic attributes, for two reasons. Firstly, because the resulting model aligns better with the domain knowledge of the practitioner. And secondly, because it is likely to lead to higher anomaly detection performance.
Finally, we note that there are other types of anomaly detection algorithms, like the Support Vector Machine (SVM) and the Isolation Forest (IF), which should also be adapted for monotonic anomaly detection.
{credits}
7.0.1 Data and code.
Datasets and the code to reproduce our experiments are available at https://liacs.leidenuniv.nl/~lenzou/code/lenz-2025-monotonic.tar.gz.
7.0.2 Acknowledgements
@@
Figure 2: Mean AUROC for NND as a function of $k$, across real-life datasets.
Appendix 0.A The number of neighbours $k$
In semi-supervised anomaly detection, we generally do not have access to anomalies for the purpose of hyperparameter tuning. Therefore, we have chosen default values in our main experiment. Nevertheless, we may ask how performance on our test sets varies as a function of different hyperparameter values. For reasons of space, we limit ourselves here to the number of neighbours $k$ of NND.
Fig. 2 displays the mean AUROC of ramp and absolute distance as a function of $k$, for the real-life datasets. We observe that the default value of $k=8$ lies close to the maximum for both distance variants, providing validation that this is a good default value.